Use the method of substitution to solve the following system of equations. If the system is dependent, express the solution sFill in the blank by performing the indicated elementary row operation(s). -4R2+R1
Answer #1

Solveds By substitution; 2x+ 5y = -21 - loxtasy = -105 from ea oj 2x+5y =-21 . 2x=- 21-sy X=-31 - Su Answer) © solved; Elemen